• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name N Toro

    Log entry time 00:09:26 on July 4,2010

    Entry number 329828

    keyword=Shift summary (Swing)

    swing shift July 3:
    SL - Natalia Toro
    TO/3rd - David Anez

    continuing alignment (led by Seamus):
    16:36 -- Carbon hole spot++ looks strange, moving to BeO viewer
    16:38 -- 5uA on BeO viewer; hard to see anything
    16:53 -- MCC is restarting fast feedback

    17:11 -- moving to Pb/D #1

    Runs (spot++)
    1567: 12x8 MCC (6x4mm) raster at (0,0) for both A and B, T4 trigger
    1568; 16x16 MCC (~8x8mm) raster at (0,0), T4 trigger
    1569: 16x16 MCC (~8x8mm) raster at (0,0), T1 trigger
    no evidence that we're hitting the frame.

    1570: 16x16 MCC, both BPMs (X=0, Y=-1), T1 trigger

    17:39: MCC is going to start ion chamber calibration for Pb/C #1 up to 30uA, with raster 12x8 in MCC units, both BPMs at (0,0).

    We'll use parasitically for temperature tests and detector checkout.

    18:30: Target moved to empty & raster off (so MCC can check beam position on dump viewer, establish baseline for heating test)

    18:39: Moving to Pb/C #1, re-establishing 12x8 MCC raster for permanent magnet heating test

    18:47 ramping to 25 uA for magnet heating test

    19:53 moving target to empty to test high-temperature beam spot

    19:57 MCC confirms beam on dump viewer, with magnet temperature of 90 degrees. No observable motion. See MCC ELOG #1532384

    20:15 Taking detector checkout runs at 0.5uA, using triggers T3,T4,T6
    (first run: 1596)

    20:24 MCC: Moller quads have been down since about 1pm, adverse effect on optics likely. They tried to cycle, but Moller quads did not come up.

    20:25 Right-HRS Q1 is down, "main power off" error.

    20:46 Requested 0.5uA on Pb/C #1, for L-HRS checkout run.

    20:54 Requesting escorted access so Alex can try to reset right-HRS Q1 power supply.

    21:30 Q1 fault was cleared.

    21:36 Request 0.5uA on Pb/C #1 for R/L-HRS checkout run.

    21:41 Moller quads are back up; checking spot++ with tune beam

    21:51 Checkout run (1600) with T3 on,

    21:54 Spot++ run (1601) on T1.

    21:56 Request controlled access to work on detectors